An entrance test is made up of three parts. Each part has 20 questions. What is the maximum number of questions one can fail and still get an overall average of 70% to pass the test?
Answer by stanbon(75887)

An entrance test is made up of three parts. Each part has 20 questions. What is the maximum number of questions one can fail and still get an overall average of 70% to pass the test?
----------
Total # of questions = 3*20 = 60.
Max number of incorrect answers = 0.30*60 = 18
=================================================
Cheers,
Stan H.
==========